OpenStudy (anonymous):

Find the sum of two numbers is 15. Twice one number equals 3 times the other. Find the numbers

OpenStudy (aroub):

x+y=15 y=15-x 2x=3y 2x=3(15-x) 2x=45-3x -45=-5x x=9 y=15-x y=15-9 y=6 The two numbers are 9 and 6.

OpenStudy (anonymous):

Thank you boo :) thanks for your help seriously

OpenStudy (aroub):

Hehe, anytime! ;D